OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. -2x+5/3y=-5 3x=7/2y=9/2 Show work PLEASE :)

OpenStudy (dumbcow):

haha, didn't we already help you solve this

OpenStudy (anonymous):

i am trying to teach my son and i need all the work i can get i just needed more info

OpenStudy (dumbcow):

ok what else did you want to know

OpenStudy (anonymous):

how to find y

OpenStudy (dumbcow):

a lot of steps were done for you in other post

OpenStudy (dumbcow):

Remember the goal when using elimination is to multiply equations by numbers or scale factors which will allow a variable to cancel out if the equations are added together

OpenStudy (anonymous):

we got y in 4 steps: 1)multiply the first equation by 3 get \[-6x+5y=-15\] 2) multiply the second equation by 2 you get \[6x-7y=9\] 3) add \[-2y=-6\] 4) divide by -2 \[y=3\]

OpenStudy (anonymous):

now you need x. replace y by 3 in either equation, get x

OpenStudy (anonymous):

he/she reposted it

OpenStudy (anonymous):

replace y by 3 \\[6x-7(3)=9\] multiply \[6x-21=9\] add 21 \[6x=30\] divide by 6 \[x=5\]